OpenStudy (anonymous):

What is total change in entropy equal to? A. The standard entropy squared multiplied by the temperature of the system B. The standard entropy squared multiplied by the number of moles C. The standard entropy multiplied by the number of moles D. The standard entropy multiplied by the temperature of the system

OpenStudy (anonymous):

A or D I think?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

If I remember correctly, \(\sf dS = \frac{ dQ}{T} \), where the change in entropy must always be positive.

OpenStudy (anonymous):

that's what I thought but that's not an option in the above answers @abb0t

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@Kainui @Abhisar

OpenStudy (kainui):

Apparently the Standard Entropy is just another name for the molar entropy. That is, the amount of entropy change of a single mole of the substance. So since standard entropy is the entropy per mole of substance, if you multiply it by the total number of mols of substance undergoing an entropy change, then you have the total entropy change. \[\LARGE \Delta S = \frac{\Delta S}{mols}*mols\]

OpenStudy (anonymous):

so the answer is C? @Kainui

OpenStudy (kainui):

Why would you think that?

OpenStudy (anonymous):

because isn't \[\frac{ \Delta S}{ mols }\] the standard entropy and then the \[\times mols\] make it entropy times mols? @Kainui

OpenStudy (anonymous):

@jdoe0001 @JoannaBlackwelder

OpenStudy (joannablackwelder):

I think @Kainui was trying to say that total change in entropy = standard change in entropy * moles

OpenStudy (joannablackwelder):

Because standard change in entropy is change in entropy per mole

OpenStudy (joannablackwelder):

ΔS=(ΔS/mols)∗mols The part in parentheses is standard entropy

OpenStudy (anonymous):

so it is C then?

OpenStudy (joannablackwelder):

That is what I get.